OpenStudy (laddiusmaximus):

find the derivative x^(3)-2x^(2)+19sqrtx +(4/x^(3))-6

OpenStudy (anonymous):

is it 3x^2 -4x +19/(2 (sqrtx)) -12/x^4

OpenStudy (laddiusmaximus):

ok i follow up to 19/2 (sqrtx) and -12/x^(4)

OpenStudy (campbell_st):

rewrite the equation \[y = x^3 -2x^2 +19x^{1/2} +4x^{-3} - 6\] \[dy/dx = 3x^2 - 4x + 19/2 x^{-1/2} -12x^{-4}\] \[dy/dx = 3x^2 - 4x + 19/(2\sqrt{x}) -12/x^4\]